OpenStudy (anonymous):

solve using completing the square leave in simplified form help plz 3x^2-6x=27

OpenStudy (anonymous):

step one, divide by 3 so the leading coefficient is 1 you will get \[x^2-2x=9\]

OpenStudy (anonymous):

step 2 take half the coefficient of the \(x\) term, which his this case is \(-1\) square it and get 1 and write \[(x-1)^2=9+1=10\]

OpenStudy (anonymous):

step 3 take the square root of both sides. don't forget the \(\pm\) to get \[x-1=\pm\sqrt{10}\]

OpenStudy (anonymous):

and the last step is to solve for \(x\) by adding one to both sides to get \[x=1\pm\sqrt{10}\]

OpenStudy (anonymous):

that sign that you made i dont know about that ive never actually learned the basics of this or the signs

OpenStudy (anonymous):

you mean \(\pm\) ?

OpenStudy (anonymous):

yeah does that just mean not equal too?

OpenStudy (anonymous):

no it means "plus or minus"

OpenStudy (anonymous):

you can write instead \(x=1+\sqrt{10}\) or \(x=1-\sqrt{10}\)

OpenStudy (anonymous):

ohhh so its a + sign and then under that itz minus

OpenStudy (anonymous):

rigjt

OpenStudy (anonymous):

i didnt know that thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

:)